# Materials and spin characteristics of amino-terminated nanodiamonds embedded with nitrogen-vacancy color centers

### v1

@@ II Methods
Our analysis (see below) revealed that the FND solutions contain sulfur, which may originate from acid cleaning residues or surfactants added to prevent FND aggregation in the solution. Furthermore, post-cleaning of the purchased FND solutions removes some organic sulfur-containing materials (e.g., surfactants or ultrasmall carbon-based particles). However, this process significantly reduces colloidal stability and impacts the examined optical properties, thus we incline to claim that sulfur-related surfactants were present in the as-received FNDs.
The FNDs were either used as is or washed before measurements using solvent exchange centrifugation with DI water three times, acetone two times, and DI water five times, in order to remove the surfactants. We called the initial samples as-received, whereas they were labeled as ”washed” when the surfactants were removed as explained above.
@@
Scanning electron microscopy (TESCAN MIRA3 SEM) and high-resolution transmission electron microscopy (Thermo Fisher Themis Cs-corrected HRTEM) were used to observe the shape and size of FNDs. Energy Dispersive Spectroscopy (EDS) spectra were recorded both in SEM and HRTEM to determine the particle composition. For this analysis, the particles were dispersed over a Si substrate in the case of EDS measurements. For the HRTEM analysis, the aqueous suspension of FNDs was drop-cast onto a TEM grid (TED Pella) covered with an ultrathin carbon layer supported by lacey carbon after the purification.
@@ III.2 Analysis of the surface termination
Comparing the FTIR spectra of as-received and washed FNDs reveals two major changes: (i) the C--H bending modes at 1460 $\text{cm}^{-1}$ Petit and Puskar (2018) decrease in intensity after washing, and (ii) the peaks around 1216 $\text{cm}^{-1}$ show a reduction. At the same time, we observed that washed FNDs, regardless of their size, lose their aqueous colloidal stability. This led us to suspect that the vibration bands near 1216 $\text{cm}^{-1}$ correspond to sulfonate groups, which contribute to the stabilization of FND particles in water. However, unambiguous identification from FTIR spectra alone remains challenging due to overlapping C--O stretching vibration bands. To test this hypothesis, we performed SEM-EDS and HRTEM-EDS elemental analyses. The results show that sulfur-related peaks are clearly visible in as-received FNDs but become negligible in the washed samples (see Figs. 8 and 9 in Appendix B). This finding suggests the presence of sulfur-containing organic molecules, or a variety of such molecules, that are weakly attached to the surface. These surfactants play a significant role in colloidal stabilization and, as discussed below, influence the optical properties of FNDs.
@@ Appendix A Sample preparation
To realize the --OH termination the following procedure was used. The FND particles were aggregated in Eppendorf centrifuge tubes then removing the supernatant Tetrahydrofuran (THF) - dried with CaCl2 - was added and the mixture was dispersed. This step removed the surfactant from the FNDs. In a round bottom flask the redispersed FND solution was diluted with 5 mL CaCl2 dried THF and added 1.6 mL of 1.0 mol LiAlH4 in THF solution. This reaction mixture was refluxed intensely overnight. After evaporating THF, the reaction mixture was quenched by adding 1 mol HCl. The cleaning and collecting took place by centrifugation using centrifuge tubes with polyethersulfone filter media. The product was washed with Millipore water several times to achieve a neutral pH condition. Finally, it was washed with acetone several times.
The amino-terminated FND particles were created by applying Hofmann-degradation. First, the FND suspension was dried. SOCl2 was added to the FNDs and sonicated at 50 C for 3 hours. The SOCl2 was removed under vacuum at 50 C and NH3/dioxane solution was added to the FNDs. This was followed by a new sequence of sonication at 50 C. The solution with NaOH and Br2 was heated at 80 C for 10 minutes. NH3 is added to the cooled mixture to remove the Br2 excess. Finally, NH3 was eliminated by evaporation in vacuum at 50 C. The pH value was set by applying HCl. The solution was cleaned by repeating the cycle of the centrifugation and suspension in water 10 times. The surfactant (since the original suspension of FND contained sulfur) was removed at the final purification step.
@@ Appendix B Analysis by SEM, HRTEM, and EDS
The SEM images clearly show the non-spherical shape of the FNDs in Fig. 7. Fig. 8 presents the EDS spectra of ND10 as-received and washed samples as obtained by SEM. The presence of sulfur is evident in the as-received samples. This observation was also proven by the EDS analysis performed in HRTEM (Fig. 9).
@@
Figure 9: HRTEM image and HRTEM-EDS analysis of (a) ND10 as-received and (b) ND10 washed particles. The scale corresponds to 10 nm. The black line indicates the measurement of the size of the nanodiamond particles.

### v2

@@ II Methods
Our analysis (see below) revealed that the FND solutions contain sulfur, which may originate from acid-cleaning residues. The FNDs were either used as is or washed before measurements using solvent exchange centrifugation with DI water three times, acetone two times, and DI water five times, in order to remove the sulfur-related residues affecting the colloid stability and the optical properties of the samples. We called the initial samples as-received, whereas they were labeled as ”washed” when the sulfur-related residues were removed as explained above.
@@ III.2 Analysis of the surface termination
Comparing the FTIR spectra of as-received and washed FNDs reveals two major changes: (i) the C--H bending modes at 1460 $\text{cm}^{-1}$ Petit and Puskar (2018) decrease in intensity after washing, and (ii) the peaks around 1216 $\text{cm}^{-1}$ show a reduction. At the same time, we observed that washed FNDs, regardless of their size, lose their aqueous colloidal stability. This led us to suspect that the vibration bands near 1216 $\text{cm}^{-1}$ correspond to sulfonate groups, which contribute to the stabilization of FND particles in water. However, unambiguous identification from FTIR spectra alone remains challenging due to overlapping C--O stretching vibration bands. To test this hypothesis, we performed SEM-EDS and HRTEM-EDS elemental analyses. The results show that sulfur-related peaks are clearly visible in as-received FNDs but become negligible in the washed samples (see Figs. 8 and 9 in Appendix B). This finding suggests the presence of sulfur-containing organic molecules, or a variety of such molecules, that are weakly attached to the surface. These sulfur-related residues play a significant role in colloidal stabilization and, as discussed below, influence the optical properties of FNDs.
@@ Appendix A Sample preparation
To realize the --OH termination the following procedure was used. The FND particles were aggregated in Eppendorf centrifuge tubes then removing the supernatant Tetrahydrofuran (THF) - dried with CaCl2 - was added and the mixture was dispersed. This step removed the sulfur-related residues from the FNDs. In a round bottom flask the redispersed FND solution was diluted with 5 mL CaCl2 dried THF and added 1.6 mL of 1.0 mol LiAlH4 in THF solution. This reaction mixture was refluxed intensely overnight. After evaporating THF, the reaction mixture was quenched by adding 1 mol HCl. The cleaning and collecting took place by centrifugation using centrifuge tubes with polyethersulfone filter media. The product was washed with Millipore water several times to achieve a neutral pH condition. Finally, it was washed with acetone several times.
The amino-terminated FND particles were created by applying Hofmann-degradation. First, the FND suspension was dried. SOCl2 was added to the FNDs and sonicated at 50 C for 3 hours. The SOCl2 was removed under vacuum at 50 C and NH3/dioxane solution was added to the FNDs. This was followed by a new sequence of sonication at 50 C. The solution with NaOH and Br2 was heated at 80 C for 10 minutes. NH3 is added to the cooled mixture to remove the Br2 excess. Finally, NH3 was eliminated by evaporation in vacuum at 50 C. The pH value was set by applying HCl. The solution was cleaned by repeating the cycle of the centrifugation and suspension in water 10 times. The sulfur-related residues were removed at the final purification step.
@@ Appendix B Analysis by SEM, HRTEM, and EDS
The SEM images clearly show the non-spherical shape of the FNDs in Fig. 7. Fig. 8 presents the EDS spectra of ND10 as-received and washed samples as obtained by SEM. The presence of sulfur is evident in the as-received samples. This observation was also proven by the EDS analysis performed in TEM (Fig. 9).
@@
Figure 9: HRTEM image and TEM-EDS analysis of (a) ND10 as-received and (b) ND10 washed particles. The scale corresponds to 10 nm. The black line indicates the measurement of the size of the nanodiamond particle.